Determine the gravitational force in \mu N

μN acting between two spheres that are touching each other. Express the result with two decimal places.

The mass of each sphere is 220 kg and the radius is 306 mm. Use as a gravitational constant value G = 66.7430 10-12 m3 kg-1 s-2

1
Expert's answer
2021-08-16T08:34:00-0400

"F=\frac{Gm^2}{(2r)^2}=\frac{Gm^2}{4r^2}=8.6~\mu N."